A voltage source 𝑣(𝑡)=240 sin(314.16𝑡−20v°)𝑉 is connected to a load having an toa load having an impedance of Z. The resulting current through the load is 𝑖(𝑡)=15 sin(314.16𝑡+22.5°)A. Determine the circuit power factor and identify the elements that constitute the load, given that Z comprises of only two elements connected in series.


1
Expert's answer
2021-06-09T04:35:03-0400

From the given values of current and voltage.

"V_{rms}=240/\sqrt2\newline
I_{rms}=15/\sqrt2\newline
Phase \space angle= 20+22.5=42.5"


This indicates the load is a capacitive load.

Magnitude of impedance"=V_{rms}\/I_{rms}"


"Z=240/15=16\varOmega\newline
R=Zcos\phi=16cos42.5=11.80\varOmega\newline
X=Zsin\phi=10.81\varOmega\newline
i.e Z=11.8 - j10. 81"


The capacitor value ="\frac{1}{X\omega}=\frac{1}{10.81\times314.16}=0.294\times10^{-3}F=0.294mF"